OpenStudy (anonymous):

Solve by elimination 5x +7y = 77 5x + 3y = 53 I got 32.5 for y but i think it is wrong can someone check plz

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@robz8 did i do this right

OpenStudy (anonymous):

subtract the equations from each other 5x + 7y = 77 -( 5x + 3y = 53 ) 4y = 24 y = 6

OpenStudy (anonymous):

i suck at this

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

just practice, it will click